OpenStudy (steven513):

The reciprocal of a number minus the reciprocal of two less than the number is -2/15. find the number.

OpenStudy (irishboy123):

\(\large \frac{1}{n} - \frac{1}{n- 2}= -\frac{2}{15}\) Yes?

OpenStudy (irishboy123):

\(n = -3, 5\)